I cant seem to get the answer for the following problem. My mother cant get it neither. Here is the problem: (2^-1*tu^-1)^3*(6tu^-1)^-2 . The section says to express each problem in simplest form, without using negative or zero exponents. Thanks, Luke
Answer by venugopalramana(3286)

Here is the problem: (2^-1*tu^-1)^3*(6tu^-1)^-2 .
I TRUST YOUR POSITIONING OF BRACKETS IS CORRECT AND AM SOLVING THE PROBLEM ACCORDINGLY.PLEASE NOTE THAT THE ANSWER IS ENTIRELY DEPENDENT ON USE OF PROPER POSITIUONING OF BRACKETS.
(2^-1*tu^-1)^3*(6tu^-1)^-2 = {{((1/2)*t*(1/U))}^3}*{{(6t*(1/U))}^-2}
={(T^3)/(8*U^3)} * {(6T/U)^-2}={(T^3)/(8*U^3)} * {(U^2)/(36T^2)}
=(T^(3-2))/(8*36*U^(3-2))=T/(288U)
The section says to express each problem in simplest form, without using negative or zero